How many liters each of a 15% antifreeze solution and a 30% antifreeze solution must be mixed to make 6 liters of a 20% antifreeze solution?
Answer by Maths68(1474)

Solution A
Amount = x
Concentration =15% =0.15
===============================================================================
Solution B
Amount = 6-x
Concentration =30% = 0.3
==============================================================================
Resultant Solution
Amount =6
Concentration =20%=0.2
===============================================================================
Solution B
[Amount Solution A * Concentration A] + [Amount Solution B * Concentration of B] = Amount of Resultant * Concentration of resultant
(x)(0.15)+(6-x)(0.3)=(6)(0.2)
0.15x+1.8-0.3x=1.2
0.15x-0.3x=1.2-1.8
-0.15x=-0.6
-0.15x/-0.15=-0.6/-0.15
x=4
=============================================================================
Solution A
Amount = x=4 liters
Concentration =15% =0.15
===============================================================================
Solution B
Amount = 6-x =6-4=2 liters
Concentration =30% = 0.3
===============================================================================
4 liters of a 15% antifreeze solution and 2 liters of 30% antifreeze solution must be mixed to make 6 liters of a 20% antifreeze solution.